Your friend is about to drive away after moving out and you notice that they left behind their basketball, so you decide to throw it at a 45LaTeX: ^{^\circ}∘ ∘ from the horizontal towards the truck. If the truck is 9.0 m away from you, at what speed must you throw the ball for it to make it into the truck? Assume you launch the ball at the same height as the truck bed.
Answer #1

given
theta = 45 degrees
Range, R = 9 m
initial speed, vo = ?

we know, horizontal range, R = vo^2*sin(2*theta)/g

==> vo^2 = R*g/sin(2*theta)

vo = sqrt(R*g/sin(2*theta))

= sqrt(9*9.8/sin(2*45))

= 9.4 m/s <<<<<<<<------------------Answer